the excess pressure inside  soap bubble id thrice the excess pressure inside a second soap bubble. what is the ratio betwen the volume of the first and the second bubble?
 

Grade:11

1 Answers

Arun
6 years ago
Dear Shimi
 
4S/r1 = 3 * 4S/r2
r2 = 3* r1
V1/V2 = (4/3)\pir1³ / (4/3)\pi r2³
V1/V2 = r 1³/ r2³ 
V1/V2 = 1/27
 
Regards
Arun (askIITians forum expert)